OpenStudy (anonymous):

PLEASE HELP EASY QUESTION Where does the graph of y=2x^2-5x+3 cross the x-axis?

OpenStudy (klimenkov):

Solve an equation \[2x^2-5x+3=0\] and you will get 2 points on the x-axis where the graph crosses it

OpenStudy (zzr0ck3r):

do you know the "AC method" ?

OpenStudy (anonymous):

No but I really need help

OpenStudy (campbell_st):

factorise the quadratic (2x - 3)(x - 1) = 0 this should help you find the intercepts

OpenStudy (anonymous):

The answer choices are -1/2,0 and -3,0 1/2,0 and 3,0 3/2,0 and 1,0 -3/2,0 and -1,0

OpenStudy (zzr0ck3r):

A*C = 6 -2*-3 = 6 and adds to -5 2x^2-5x+3= 2x^2 - 3x - 2x + 3 = x(2x-3)-1(2x-3) = (2x-3)((x-1) set equal to 0 (2x-3)(x-1)=0 2x-3=0 or x-1=0 2x=3 or x=1 x=2/3 or x = 1

OpenStudy (zzr0ck3r):

there are two other ways to do this, one being "complete the square" and another is with the equation (-b +- sqrt(b^2-4ac))/2a but the method I showed is the first way I learned and its easy if your not used to completing the square